OpenStudy (anonymous):

03. What are the first four terms of the sequence represented by the expression n(n – 1) – 4? (1 point) –2, 2, 6, 10 –3, 0, 3, 6 –4, –2, 2, 8 –5, –4, –2, 6

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

did they give you a of 1?

OpenStudy (anonymous):

no.. that was all the information

OpenStudy (anonymous):

@awkwardpanda do you know this???

OpenStudy (anonymous):

your difference is -4 so which sequence decreases by 4 every time?

OpenStudy (anonymous):

None of them decrease.

OpenStudy (anonymous):

or wait....are you sure its not d(n-1)-4?

OpenStudy (anonymous):

please double check for that because that will make a huge difference

OpenStudy (anonymous):

Yes. Im 100% sure.

OpenStudy (anonymous):

hmm...alright. ill keep trying to figure it out. sorry for the wait

OpenStudy (anonymous):

its cool

OpenStudy (anonymous):

you can solve for a of 1. making n=1 so you have 1(1-1)-4 = -4 as your first term

OpenStudy (anonymous):

then if you do a of 2. making n=2 you get 2(2-1)-4 so then 2-4= -2

OpenStudy (anonymous):

then a of 3, n = 3. 3(3-1)-4. 6-4 = 2

OpenStudy (anonymous):

a of 4, n=4 4(4-1) - 4 = 12-4 = 8

OpenStudy (anonymous):

that should be your answer. wow....haha it shouldnt have taken me this long to think that out XD

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

sure thing :)